# Multi-Parameter Rating Methodology for Systematic Comparison of Theoretical Models with Experimental Data in Heavy-Ion Physics

### v1

@@ Abstract
We introduce a novel multi-parameter rating methodology for objective comparison of theoretical models with experimental data in heavy-ion collisions, addressing fundamental limitations of the traditional global $\chi^{2}$/ndf criterion. The methodology divides phase space into seven physically motivated kinematic regions reflecting distinct production mechanisms from thermal freeze-out to perturbative QCD. Each region receives a quality score $Q_{i}\in[10,1000]$ via logarithmic transformation of local $\chi^{2}_{i}/\nu_{i}$ statistics, ranging from 10 (very poor agreement) to 1000 (perfect agreement). A composite rating $R$ aggregates weighted average, geometric mean, and minimum scores with dispersion penalty, preventing compensation of poor agreement in one region by artificial success elsewhere. Demonstrated on real ALICE and representative LHCb data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons at $\sqrt{s_{NN}}=5.02$ TeV in p-Pb collisions, the methodology reveals that no universal model exists: PYTHIA8 excels for mesons (R=879) via nuclear PDFs, while models with coalescence are potentially superior for baryons. The baryon anomaly peak serves as critical discriminator, with highest quality scores $Q_{i}\sim 950$–1000 in zones 3–4. The methodology is transparent, reproducible, fully algorithmic, and ready for integration into standard analysis frameworks.
@@ 1 Introduction
The theoretical description of heavy-ion collisions at ultrarelativistic energies remains one of the most challenging problems in modern nuclear physics. Multiple models—ranging from hydrodynamic approaches (EPOS-LHC) to microscopic transport (AMPT, PHSD) to perturbative QCD generators (PYTHIA8, HIJING)—attempt to capture the complex interplay of initial-state nuclear effects, quark-gluon plasma (QGP) formation, collective flow, and hadronization [1, 2, 3].
@@
$\chi^{2}_{\text{global}}=\sum_{i=1}^{N}\frac{(y_{i}^{\text{data}}-y_{i}^{\text{model}})^{2}}{\sigma_{i}^{2}},\quad\chi^{2}/\text{ndf}$ (1)
where $N$ is the total number of data points across all kinematic regions, and ndf =Nk=N-k accounts for $k$ model parameters. While statistically rigorous, this approach suffers from a fundamental flaw: it masks critically important physical information about regional model performance.
@@ 1.2 Goal and Key Innovation
    Statistical rigor: Preserving correctness of least-squares method
@@
    Physical motivation: Space division reflects real production mechanisms
@@
    Transparency: Results are intuitively understandable
@@
    Reproducibility: Complete algorithmization
@@
    Scalability: Generalization to arbitrary number of models and observables
@@
1.3 Structure
Section 2 presents the complete methodology: zone division with physical justification, quality score transformation, weight coefficients, aggregation formula, visualization tools, and database structure. Section 3 demonstrates application to real ALICE and representative LHCb data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons in p-Pb collisions at 5.02 TeV, revealing forward/backward asymmetries, baryon anomaly via coalescence, and model complementarity. Section 4 provides final model rankings. Section 4.1 presents detailed calculation example for PYTHIA8. Sections 5–7 discuss advantages over traditional $\chi^{2}$/ndf, development perspectives, and conclusions.
@@ 2.1 Why Exactly Seven $p_{T}$ Zones?
The choice of exactly seven kinematic zones is based on qualitative changes in dominant particle production mechanisms across the transverse momentum spectrum. This division has been validated through extensive studies of heavy-ion collisions at RHIC and LHC [1, 4, 5].
@@
1 <0.8<0.8 Thermal production. Boltzmann-Gibbs spectrum with chemical freeze-out temperature $T_{\text{ch}}\approx 156$ MeV. Dominated by hydrodynamic fireball expansion [2].
@@
2 0.80.81.61.6 Hydrodynamic flow. Radial flow becomes visible. Elliptic flow $v_{2}$ starts rising. Mass ordering effects appear: heavier particles acquire larger $\langle p_{T}\rangle$ [1].
@@
3 1.61.62.52.5 Cronin effect. Enhancement through initial-state multiple soft parton scattering in nuclear collisions. For baryons: onset of baryon anomaly via coalescence [6].
@@
4 2.52.54.04.0 Transition to fragmentation. Critical zone! Largest spread between models. Baryon anomaly peaks ($\Lambda/K^{0}_{S}\sim 0.95$). Interplay of soft collective dynamics and hard partonic processes. Most sensitive to QGP signatures [5].
@@
5 4.04.06.06.0 Hard fragmentation. Perturbative QCD processes dominate. Jet quenching visible in Pb–Pb. Baryon anomaly disappears: $\Lambda/K^{0}_{S}\to 0.5$ (vacuum fragmentation baseline) [4].
@@
6 6.06.010.010.0 Jet region. Pure fragmentation from hard-scattered partons. Factorized pQCD regime. Nuclear modification factor $R_{AA}$ plateaus [4].
@@
7 >10.0>10.0 Asymptotic regime. Universal fragmentation functions independent of collision system. Baseline for normalization and systematic uncertainties [7].
@@
Table 1 summarizes the physical motivation. Why exactly seven? Fewer zones (3–4) cannot separate qualitatively different regimes: thermal production, flow, Cronin enhancement, transition, hard jets, and asymptotic pQCD would be conflated. More zones (10–15) lead to low statistics per zone ($\nu_{i}<5$), causing $\chi^{2}_{i}/\nu_{i}$ fluctuations that destroy statistical reliability.
Seven zones provide optimal balance: $\nu_{i}\approx 5$–15 points ensure reliable $\chi^{2}_{i}$ calculation, clear separation of physical regimes, and convenient visualization (heptagonal spider plots). Stability tests show that $\pm 10\%$ shifts in zone boundaries change quality scores $Q_{i}$ by less than 15%, confirming robustness.
@@
Note on high-$p_{T}$ data (zones 5–7): Direct experimental measurements of particle ratios ($K^{0}_{S}/\pi$, $\Lambda/\pi$) in p-Pb collisions are currently limited to $p_{T}\lesssim 4.5$ GeV/c [8, 9]. For zones 5–7 ($p_{T}>4.5$ GeV/c), we apply a theoretically motivated approximation based on Pb-Pb data from ALICE [8, 10]. This approach is justified by two complementary theoretical arguments: (i) universality of particle ratios at high multiplicity demonstrated by Shao et al. [11], who showed remarkable similarity in yield ratios (including $K/\pi$, $\Lambda/\pi$) between p-Pb and Pb-Pb systems at high event multiplicity, indicating common underlying hadronization mechanisms; (ii) saturation of collective effects at high $p_{T}$ as evidenced by Chatterjee et al. [12], who observed convergence of elliptic flow patterns for particles with $p_{T}>8$ GeV/c in p-Pb and Pb-Pb collisions, suggesting that hard partonic processes dominate over system-size effects in this regime. The threshold $p_{T}=4.5$ GeV/c marks the onset of this transition region where perturbative QCD fragmentation begins to overwhelm collective soft dynamics. While this approximation introduces additional systematic uncertainty (estimated at 10–20% based on known differences in nuclear modification factors $R_{pPb}$ vs $R_{AA}$), it enables full 7-zone methodology coverage pending direct p-Pb measurements at higher $p_{T}$. All figures clearly indicate which zones use experimental data versus extrapolation.
@@ 2.2 Local Statistics and Quality Score Scale
  • (a)
    Logarithmic dependence: Transition $R_{i}\to 2R_{i}$ produces approximately constant $\Delta Q_{i}$ regardless of absolute scale, reflecting that relative deterioration matters more than absolute level.
  • (b)
@@ 2.3 Weight Coefficients
The sum $\sum_{i=1}^{7}w_{i}=10.00$ normalizes the weighted average. Zone 4 receives the highest weight ($w_{4}=2.94$) as the transition region where soft and hard processes compete, making it maximally sensitive to model physics. For asymmetric collision systems (p-Pb), a configuration modifier is applied: $w_{i,\text{Backward}}=1.2\times w_{i}$ to emphasize the backward direction where nuclear shadowing effects are strongest.

### v2

# A Rating Quality Methodology for the Theoretical Description of Experimental Data

@@ Abstract
We introduce a novel multi-parameter rating methodology for objective comparison of theoretical models with experimental data in heavy-ion collisions, addressing fundamental limitations of the traditional global $\chi^{2}$/ndf criterion. The methodology divides phase space into seven physically motivated kinematic regions reflecting distinct production mechanisms from thermal freeze-out to perturbative QCD. Each region receives a quality score $Q_{i}\in[10,1000]$ via logarithmic transformation of local $\chi^{2}_{i}/\nu_{i}$ statistics. A composite rating $R$ aggregates weighted average, geometric mean, and minimum scores with a bounded dispersion penalty, preventing compensation of poor agreement in one region by artificial success elsewhere. The seven-zone division is physically motivated and independently validated through boundary significance tests on CMS PbPb data at $\sqrt{s_{NN}}=5.02$ TeV: four of six physical boundaries are confirmed statistically significant ($p<0.05$) while none of the data-driven K=9K{=}9 candidate boundaries pass the test. Coefficient sensitivity analysis confirms complete ranking stability: ${\pm}20\%$ variations in aggregation coefficients produce $\Delta R<2\%$ with zero rank changes. Demonstrated on real ALICE data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons at $\sqrt{s_{NN}}=5.02$ TeV in p-Pb collisions, and on synthetic benchmark data consistent with LHCb kinematic coverage for the $\Lambda/K^{0}_{S}$ ratio, the methodology reveals that no universal model exists: PYTHIA8 excels for mesons ($R=879$) via nuclear PDFs, while models incorporating coalescence mechanisms are superior for baryons, as quantified by the PHSD base rating $R=893$. The baryon anomaly peak serves as the critical discriminator, with quality scores $Q_{i}\sim 950$$1000$ in zones 3–4 for coalescence models versus $Q_{i}\sim 400$$600$ for pure fragmentation generators. The methodology is transparent, reproducible, fully algorithmic, and ready for integration into standard analysis frameworks.
Keywords:
heavy-ion collisions, model comparison, chi-squared statistics, nuclear modification factor, baryon anomaly, $K^{0}_{S}$, $\Lambda$ hyperon, p-Pb collisions, ALICE, LHCb
@@ 1 Introduction
The theoretical description of heavy-ion collisions at ultrarelativistic energies remains one of the most challenging problems in modern nuclear physics. Multiple models — ranging from hydrodynamic approaches (EPOS-LHC) to microscopic transport (AMPT, PHSD) to perturbative QCD generators (PYTHIA8, HIJING) — attempt to capture the complex interplay of initial-state nuclear effects, quark-gluon plasma (QGP) formation, collective flow, and hadronization alice2024 ; heinz2013 ; bernhard2016 .
@@
$\chi^{2}_{\text{global}}=\sum_{i=1}^{N}\frac{(y_{i}^{\text{data}}-y_{i}^{\text{model}})^{2}}{\sigma_{i}^{2}},\qquad\chi^{2}/\text{ndf}$ (1)
where $N$ is the total number of data points across all kinematic regions, and $\text{ndf}=N-k$ accounts for $k$ model parameters. While statistically rigorous, this approach suffers from a fundamental flaw: it masks critically important physical information about regional model performance.
@@ 1.2 Goal and Key Innovation
    Statistical rigor: Preserving correctness of least-squares method.
@@
    Physical motivation: Space division reflects real production mechanisms.
@@
    Transparency: Results are intuitively understandable.
@@
    Reproducibility: Complete algorithmization.
@@
    Scalability: Generalization to arbitrary number of models and observables.
@@
1.3 Structure of the Paper
Section 2 presents the complete methodology: zone division with physical justification, quality score transformation, weight coefficients, aggregation formula, visualization tools, and database structure. Section 3 demonstrates application to real ALICE data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons in p-Pb collisions at 5.02 TeV, and to a synthetic benchmark for the $\Lambda/K^{0}_{S}$ ratio consistent with LHCb kinematics, revealing forward/backward asymmetries, baryon anomaly via coalescence, and model complementarity. Section 4 provides final model rankings. Section 4.1 presents a detailed calculation example for PYTHIA8. Sections 57 discuss advantages over traditional $\chi^{2}$/ndf, development perspectives, and conclusions.
@@ 2.1 Why Exactly Seven $p_{T}$ Zones?
The choice of exactly seven kinematic zones is based on qualitative changes in dominant particle production mechanisms across the transverse momentum spectrum. This division has been validated through extensive studies of heavy-ion collisions at RHIC and LHC alice2024 ; cms2017 ; jet2020 .
@@
1 <0.8<0.8 Thermal production. Boltzmann-Gibbs spectrum with chemical freeze-out temperature $T_{\text{ch}}\approx 156$ MeV. Dominated by hydrodynamic fireball expansion heinz2013 .
@@
2 0.80.81.61.6 Hydrodynamic flow. Radial flow becomes visible. Elliptic flow $v_{2}$ starts rising. Mass ordering effects appear: heavier particles acquire larger $\langle p_{T}\rangle$ alice2024 .
@@
3 1.61.62.52.5 Cronin effect. Enhancement through initial-state multiple soft parton scattering in nuclear collisions. For baryons: onset of baryon anomaly via coalescence phsd2012 .
@@
4 2.52.54.04.0 Transition to fragmentation. Critical zone! Largest spread between models. Baryon anomaly peaks ($\Lambda/K^{0}_{S}\sim 0.95$). Interplay of soft collective dynamics and hard partonic processes. Most sensitive to QGP signatures jet2020 .
@@
5 4.04.06.06.0 Hard fragmentation. Perturbative QCD processes dominate. Jet quenching visible in Pb–Pb. Baryon anomaly disappears: $\Lambda/K^{0}_{S}\to 0.5$ (vacuum fragmentation baseline) cms2017 .
@@
6 6.06.010.010.0 Jet region. Pure fragmentation from hard-scattered partons. Factorized pQCD regime. Nuclear modification factor $R_{AA}$ plateaus cms2017 .
@@
7 >10.0>10.0 Asymptotic regime. Universal fragmentation functions independent of collision system. Baseline for normalization and systematic uncertainties pythia8 .
@@
Table 1 summarizes the physical motivation. Why exactly seven? Fewer zones (3–4) cannot separate qualitatively different regimes: thermal production, flow, Cronin enhancement, transition, hard jets, and asymptotic pQCD would be conflated. More zones (10–15) lead to low statistics per zone ($\nu_{i}<5$), causing $\chi^{2}_{i}/\nu_{i}$ fluctuations that destroy statistical reliability. Seven zones provide optimal balance: $\nu_{i}\approx 5$–15 points ensure reliable $\chi^{2}_{i}$ calculation, clear separation of physical regimes, and convenient visualization (heptagonal spider plots). Stability tests show that $\pm 10\%$ shifts in zone boundaries change quality scores $Q_{i}$ by less than 15%, confirming robustness.
Choice of zone-division strategy. Three methodological approaches exist for defining zone boundaries: (i) data-driven, where $K$ is selected by minimizing information criteria (AIC/BIC) on the specific dataset; (ii) hybrid, combining a physics-motivated boundary pool with statistical optimization; and (iii) physically motivated, where each boundary corresponds to a documented qualitative change in the dominant production mechanism. We adopt approach (iii) for three independent reasons. First, AIC/BIC optimizes boundary placement for the specific dataset at hand, so results are not reproducible across experiments with different $p_{T}$ coverage, binning, or observable: changing the pool of candidate thresholds shifts the statistical optimum, making the zone division non-unique. Second, boundary significance tests (section 2.2) confirm that 4 of 6 physical boundaries are statistically significant on CMS data ($p<0.05$), while none of the data-driven K=9K{=}9 candidates pass the test independently. Third, physically motivated boundaries are defined independently of the observable being analyzed, enabling direct comparison across $K^{0}_{S}$, $\Lambda$, $D^{0}$, and flow harmonics without rebinning — a prerequisite for the multi-observable extension discussed in section 6.
Why data-driven selection degenerates to reweighted $\chi^{2}$/ndf. AIC/BIC minimization places zone boundaries where the observable’s gradient is steepest — mathematically, it maximizes inter-zone $\chi^{2}$ across candidate thresholds. Model evaluation within such zones still uses local $\chi^{2}_{i}/\nu_{i}$, so the combined procedure yields an overall metric that reduces to a dataset-specific reweighted $\chi^{2}$/ndf: zone boundaries are placed where the data changes most rapidly, zones implicitly receive weights proportional to local data density, and the physical hierarchy of production mechanisms is erased. Data-driven zoning thus recovers the very averaging problem it was designed to solve, adding computational complexity without physical gain. A concrete illustration: in the $K^{0}_{S}$ dataset of table 8, zone 2 contains $N_{2}=15$ points (most populated) while zone 4 contains $N_{4}=8$. AIC/BIC therefore assigns higher implicit weight to zone 2 — the hydrodynamic flow regime — at the expense of zone 4, the QGP-sensitive transition region assigned the highest physical weight $w_{4}=2.94$ in our scheme. Data-driven zoning thus inverts the physically motivated hierarchy precisely where model discrimination matters most.
Cross-system incommensurability of data-driven zones. Optimal data-driven boundaries are dataset-specific: the K=9K{=}9 partition of ALICE $K^{0}_{S}$ spectra in p-Pb differs from the partition of CMS $K^{0}_{S}$ spectra in Pb-Pb, since jet quenching reshapes the high-$p_{T}$ spectrum, shifting gradient maxima. Consequently, model ratings computed with data-driven zones are incommensurable across experiments and collision systems — a model’s score on ALICE p-Pb data cannot be directly compared with its score on CMS Pb-Pb data, defeating the purpose of a community-wide evaluation framework. In contrast, physically motivated boundaries encode universal QCD properties: the chemical freeze-out temperature $T_{\text{ch}}\approx 156$ MeV heinz2013 , the onset of radial flow, the Cronin enhancement threshold, and the soft-to-hard transition scale are all set by $\Lambda_{\text{QCD}}$ and are system-independent. The same zone structure $\{0.8,\,1.6,\,2.5,\,4.0,\,6.0,\,10.0\}$ GeV/$c$ applies directly to p-Pb, Pb-Pb, Xe-Xe, and future O-O collisions, enabling genuinely comparative model rankings across the full LHC heavy-ion programme.
Why the hybrid approach does not resolve the problem. The hybrid scheme — a fixed physics-motivated candidate pool with statistical optimization — preserves reproducibility only if the candidate pool is defined a priori and is identical across all analyses. Statistical optimization then selects from this pool based on inter-zone $\chi^{2}$ differences in the specific dataset, reintroducing data dependence. More fundamentally, if physical boundaries carry their meaning from the documented change in dominant production mechanism, then adjusting their positions to fit the current dataset destroys the independence between zone definition and model evaluation: the same data that defines the zones is used to score models within them, introducing a circularity absent in the purely physical scheme.
Adaptive Zone Merging Rule. When $\nu_{i}<3$, adjacent high-$p_{T}$ zones are merged. If $N_{7}<4$, zones 6 and 7 are combined into $6.0\infty$] with weight $w_{6+7}=1.77$ and the analysis proceeds with $K_{\rm eff}=6$. For the p-Pb configurations analysed here ($N_{7}^{\rm Bwd}=2$, $N_{7}^{\rm Fwd}=3$), zones 6+7 are merged in both directions; tables 9 and 10 therefore use $K_{\rm eff}=6$ for the p-Pb system, while full $K=7$ applies for Pb–Pb where $N_{7}\geq 5$.
2.2 Boundary Significance Validation
To validate the $K=7$ boundary choice independently from data, we apply Welch $t$-tests and Likelihood Ratio Tests (LRT) using digitized CMS PbPb data at $\sqrt{s_{NN}}=5.02$ TeV, 0–5% centrality cms2017 ($N=27$, 0.750.751414 GeV/$c$). The null hypothesis is $H_{0}:\langle R_{AA}\rangle_{\rm left}=\langle R_{AA}\rangle_{\rm right}$. Although the validation uses Pb–Pb data, the physical thresholds being tested — onset of Cronin enhancement, soft-to-hard transition, and high-$p_{T}$ pQCD plateau — are properties of QCD itself and manifest in both AA and pA systems, as confirmed by comparative studies cms2017 ; alice2024 .
Table 2: Boundary significance tests on CMS PbPb 5.02 TeV data cms2017 . Physical boundaries: 4/6 confirmed ($p<0.05$). Data-driven K=9K{=}9 candidates: 0/3 confirmed. The K=9K{=}9 boundary at 3.0 GeV/$c$ is an artifact: any threshold placed within [2.5–4.0] bisects the steep $R_{AA}$ descent already captured by the physical boundary at 2.5 GeV/$c$ and therefore yields $p<0.001$ without carrying independent physical information. $R_{AA}$ plateau ($\approx 0.13$) on both sides; boundary independently motivated by $\Lambda/K^{0}_{S}$ and $v_{2}$ observables.
Boundary (GeV/$c$) Type $p$ Result
$0.8$ Physical no data left of boundary
$1.0$ K=9K{=}9 0.491 NS
$1.6$ Physical 0.003 $p<0.01$
$2.5$ Physical 0.0002 $p<0.001$
$3.0$ K=9K{=}9 0.0001 artifact
$4.0$ Physical 0.003 $p<0.01$
$6.0$ Physical 0.910 NS
$8.0$ K=9K{=}9 0.814 NS
$10.0$ Physical 0.038 $p<0.05$
Physical K=7K{=}7 boundaries are supported by independent statistical tests. The three K=9K{=}9 data-driven boundaries carry no independent significance.
@@ 2.2 Boundary Significance Validation
Note on high-$p_{T}$ data (zones 5–7): Direct experimental measurements of particle ratios ($K^{0}_{S}/\pi$, $\Lambda/\pi$) in p-Pb collisions are currently limited to $p_{T}\lesssim 4.5$ GeV/$c$ alice2013prl ; lhcb2018 . For zones 5–7 ($p_{T}>4.5$ GeV/$c$), we apply a theoretically motivated approximation based on Pb-Pb data from ALICE alice2013prl ; alice2014plb . This approach is justified by two complementary theoretical arguments: (i) universality of particle ratios at high multiplicity demonstrated by Shao et al. shao2017prc , who showed remarkable similarity in yield ratios (including $K/\pi$, $\Lambda/\pi$) between p-Pb and Pb-Pb systems at high event multiplicity, indicating common underlying hadronization mechanisms; (ii) saturation of collective effects at high $p_{T}$ as evidenced by Chatterjee et al. chatterjee2025prl , who observed convergence of elliptic flow patterns for particles with $p_{T}>8$ GeV/$c$ in p-Pb and Pb-Pb collisions, suggesting that hard partonic processes dominate over system-size effects in this regime. The threshold $p_{T}=4.5$ GeV/$c$ marks the onset of this transition region where perturbative QCD fragmentation begins to overwhelm collective soft dynamics. While this approximation introduces additional systematic uncertainty (estimated at 10–20% based on known differences in nuclear modification factors $R_{pPb}$ vs $R_{AA}$), it enables full 7-zone methodology coverage pending direct p-Pb measurements at higher $p_{T}$. All figures clearly indicate which zones use experimental data versus extrapolation.
@@ 2.3 Local Statistics and Quality Score Scale
  1. (a)
    Logarithmic dependence: The transition $R_{i}\to 2R_{i}$ produces approximately constant $\Delta Q_{i}$ regardless of absolute scale, reflecting that relative deterioration matters more than absolute level.
  2. (b)
@@ 2.4 Weight Coefficients
The weight $w_{4}=2.94$ (the largest) encodes the physical prior that QGP signatures are maximally discriminating in zone 4: here, soft collective and hard perturbative mechanisms compete most strongly, the baryon anomaly peaks, and $R_{AA}$ undergoes its steepest descent — all direct QGP diagnostics. A data-driven weighting would assign weights proportional to data density (table 8: $N_{2}=15>N_{4}=8$), promoting the hydrodynamic-flow zone 2 over the QGP-sensitive transition zone 4 and precisely inverting the physical hierarchy. The physically motivated weights thus encode domain knowledge that is irreducible to statistical counting.
@@
The sum $\sum_{i=1}^{7}w_{i}=10.00$ normalizes the weighted average. Zone 4 receives the highest weight ($w_{4}=2.94$) as the transition region where soft and hard processes compete, making it maximally sensitive to model physics. For asymmetric collision systems (p-Pb), a configuration modifier $w_{i,\text{Backward}}=1.2\times w_{i}$ is applied to emphasize the backward direction where nuclear shadowing effects are strongest. Since $Q_{\text{weighted}}$ is computed as $\sum w_{i}Q_{i}/\sum w_{i}$, a uniform multiplicative modifier cancels in the ratio and does not alter $Q_{\text{weighted}}$ itself; it does, however, affect $\sigma(Q)$ and $Q_{\text{geometric}}$ through zone-specific weighting, producing a modest enhancement of backward-direction sensitivity in those two aggregators.